Given a normal distribution with u = 100 and o = 10,

c. What is the probability that X < 75 or X>110? The probability that X < 75 or X>110 is .8475.
Answer #1

Ans- 100, 8= 10 - XN Normal (4,8) We want to find, P(x275 os x) 110) = P(x(75) + P(x > 110) = P ( x 754) + P(x10 ** (z ( 75-